The long leg of a right triangle is 17 meters longer than the short leg. The hypotenuse is 25 meters long. Find the length of the short leg.
Answer by mananth(16949)

short leg-----x
long leg --------------x+17
hypotenuse--------------25
By Pythagoras theorem
x^2+(x+17)^2= 25^2
x^2+x^2+34x+289=625
2x^2+34x-336=0
/2
x^2+17x-168=0
x^2+24x-7x-168=0
x(x+24)-7(x+24)=0
(x+24)(x-7)=0
x=7 is the short leg
7+17 = 24 is the long leg
7,24,25